OpenStudy (anonymous):

Find A when 0º≤A≤90º and 3tan2A = 2tanA + 1. (Enter only the number.)

OpenStudy (dumbcow):

\[\tan (2A) = \frac{2\tan A}{1-\tan^{2} A}\] subbing that in \[6 \tan A = (2\tan A+1)(1-\tan^{2}A)\] \[2\tan^{3} A+\tan^{2}A+4\tan A-1 = 0\] now you have to solve the cubic \[2x^{3}+x^{2}+4x-1 = 0\] http://www.wolframalpha.com/input/?i=2x%5E3%2Bx%5E2%2B4x-1+%3D+0 \[x = 0.23058\] \[\tan A = 0.23058\] \[A = \tan^{-1} (0.23058)\]

OpenStudy (dumbcow):

this seems more complicated than it should be...can you clarify i read your equation correctly? is that tan^2 or tan(2A) ??